Determine if the improper integral La de converges. If it does converge, find its 22 value.
Answer #1

Solution: = [<*dx -11 -2+1 x -2+1 Compute the boundaries lim X0+ $(x)dx= F(b)-F (a) = lim (F(x))– lim (F(x)) 0 :) - lim 80+ =